# On How Iterative Magnitude Pruning Discovers Local Receptive Fields in Fully Connected Neural Networks

### v1

@@ Appendix A PCA and ICA Experiments
sklearn.decomposition.PCA and sklearn.decomposition.FastICA, setting the number of components to $n=64$. PCA and ICA were applied to all $50,000$ validation images in the ImageNet32 dataset. The results are shown in Fig. 2.
Appendix B ImageNet32 Experiments
B.1 ImageNet32-GP
@@
To construct the Gaussian clone of ImageNet32 [49], we utilize the same approach previously used to create clones of CIFAR-10 [26]. Namely, we sampled from 2-D Gaussian distributions that were fit to 100,000 images from the ImageNet32 dataset. Each color channel and image class had its own fit. This led to an average of $100$ images per class used for the fit. While not a small number, we recognize that this is a limitation. With the 2D Gaussian fits, we sampled the same number of images as the original ImageNet32: 1,281,167 train images and 50,000 test images. Because only $\approx 8\%$ of the images were used to generate the model, the distribution of number images per class is not guaranteed to match the true distribution. Because we use the Gaussian clone to assess the emergence of local RFs, and not the performance capability of the FCN models on ImageNet32-GP, we do not believe this difference affects our conclusion.
B.2 FCN model
@@
For our FCN experiments, we follow the approach of Pellegrini and Biroli (2022) [17]. We use the official implementation publicly available [^4] https://github.com/phiandark/SiftingFeatures, with three minor modifications to reduce computational costs. First, we scale the FCN down to have two hidden layers, instead of three. Given that Pellegrini and Biroli (2022) [17] found similar localization of IMP masks when using FCNs of greater widths, we reasoned that the width of the FCN does not play a major role in the emergence of local RFs. Our results (Fig. 1) demonstrate that IMP finds local RFs, even in this shallower FCN architecture. Second, we decrease the batch size from $1000$ to $40$. Given that a larger learning rate to batch size ratio has been found to lead SGD to converge to more generalizable solutions [60], we reasoned that – if anything – this choice would lead to greater localization (as local RFs are a generalizable computation). Lastly, we decreased the total number of training iterations from $100,000$ to $40,000$. Given that Pellegrini and Biroli (2022) [17] found similar (but slightly weaker) results when training for $10,000$ training steps, we reasoned that $40,000$ training steps would be a good intermediary. For details on all other architecture and training hyper-parameters, see Table S1.

### v3

@@ Appendix B PCA and ICA Experiments
We perform principal component analysis (PCA) and independent component analysis (ICA) [52] using sklearn.decomposition.PCA and sklearn.decomposition.FastICA, setting the number of components to $n=64$. PCA and ICA were applied to all $50,000$ test images in the ImageNet32 dataset. The results are shown in Fig. 2.
Given that the IMP masks and ICA components are localized, we investigate whether they overlap in the features they extract. To do this, we compute the cosine similarity between every ICA component and each IMP mask. We denote the largest cosine similarity, across all 64 ICA components, as the “cosine similarity” between the IMP mask and the ICA components. Example matches (with the mean cosine similarity scores) are shown in Fig. S2A. We note that this approach is limited. For instance, there may be transformations (e.g., rotations) that could make the IMP masks more aligned to the ICA components. However, we believe this is still a valuable approximation of the similarity between the ICA and IMP. Overall, the distribution of cosine similarity is skewed towards small values (Fig. S2B), indicating that ICA and IMP extract largely distinct features.
Figure S2: IMP and ICA find different local features. (A) Example IMP masks (top row) and the ICA component they have the highest cosine similarity with (bottom row). Masks are from three independent seeds, each chosen as having the mean cosine similarity across the population of masks considered. (B) Distribution of cosine similarity between IMP and ICA across three independent seeds.
Appendix C ImageNet32 Experiments
C.1 ImageNet32-GP
@@
To construct the Gaussian clone of ImageNet32 [53], we utilize the same approach previously used to create clones of CIFAR-10 [28]. Namely, we sampled from two-dimensional Gaussian distributions that were fit to 100,000 images from the ImageNet32 dataset. Each color channel and image class had its own fit. This led to an average of $100$ images per class used for the fit. While not a small number, we recognize that this is a limitation. With the two-dimensional Gaussian fits, we sampled the same number of images as the original ImageNet32: 1,281,167 train images and 50,000 test images. Because only $\approx 8\%$ of the images were used to generate the model, the distribution of number images per class is not guaranteed to match the true distribution. Because we use the Gaussian clone to assess the emergence of local RFs, and not the performance capability of the FCN models on ImageNet32-GP, we do not believe this difference affects our primary conclusion.
C.2 FCN model
@@
For our FCN experiments, we follow the approach of Pellegrini and Biroli [17]. We use the official implementation publicly available [^3] https://github.com/phiandark/SiftingFeatures, with three minor modifications to reduce computational costs. First, we scale the FCN down to have two hidden layers, instead of three. Given that Pellegrini and Biroli [17] found similar localization of IMP masks when using FCNs of greater widths, we reasoned that the width of the FCN does not play a major role in the emergence of local RFs. Our results (Fig. 1) demonstrate that IMP finds local RFs, even in this shallower FCN architecture. Second, we decrease the batch size from $1000$ to $40$. Given that a larger learning rate to batch size ratio has been found to lead SGD to converge to more generalizable solutions [63], we reasoned that – if anything – this choice would lead to greater localization (as local RFs are a generalizable computation). Lastly, we decreased the total number of training iterations from $100,000$ to $40,000$. Given that Pellegrini and Biroli [17] found similar (but slightly weaker) results when training for $10,000$ training steps, we reasoned that $40,000$ training steps would be a good intermediary. For details on all other architecture and training hyper-parameters, see Table S1.
@@
Appendix D Second layer preactivation kurtosis
To test whether IMP increases the non-Gaussianity of later FCN layers, we compute the preactivation kurtosis of the second hidden layer, as a function of IMP round (as noted in Appendix C.2, our FCN architecture only contains two hidden layers). Note that in this case, the preactivation equation (Eq. 3) is given by
$\lambda_{i}(X)=\sum_{j=1}^{N}W_{ij}\tilde{a}_{j}(X)+b_{i},$ (7)
where $\tilde{a}_{j}(X)$ is the activation of the $j^{\text{th}}$ hidden unit in the first hidden layer (Eq. 2), with batch normalization. That is, $\tilde{a}_{j}(X)=\gamma_{j}a_{j}(X)+\beta_{j}$, where $\gamma_{j}$ and $\beta_{j}$ are the learned parameters for the second layer.
As with the preactivations of the first hidden layer, we find that kurtosis increases with rounds of IMP (Fig. S3), plateauing earlier than the preactivation kurtosis of the first layer (Fig. 4B). The preactivation kurtosis slightly decreases at higher rounds of IMP (Fig. S3). This difference may be due to the fact that the second layer units are getting inputs from the hidden units in the first layer, as opposed to inputs from the images. Future work can explore the localization in later layers, and how it compares to localization in the first hidden layer, in more detail.
Figure S3: Preactivation kurtosis increases with IMP. Same as Fig. 4B, but when computing the preactivation kurtosis of the FCN’s second layer.
Figure S4: Cavity score is not strongly connected to parameter value (A)–(B) Parameter value, as a function of cavity score, for IMP round 0 and 9. Each dot corresponds to one non-pruned weight.